Question 8 Assuming an R0=7.7 for Measles, what percentage of the population needs to be vaccinated to achieve herd immunity? Round to the nearest whole percentage. V= (1-(1/7.7))*100=______
Answer #1

V= (1-(1/7.7))*100

V = (1 - 0.12987) * 100

V = 0.87013 * 100

V = 87.013

ie V = 87 %